OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, -4) with slope 4. A. y + 4 = 4 (x+4) B. y + 4 = 4 (x - 4) C. y - 4 = 4 (x+ 4)

Nnesha (nnesha):

mind posting one question at a time please ???

OpenStudy (anonymous):

You have your coordinates:\[\large (x_1,y_1)\] Point-Slope Formula:\[\large y - y _{1}= m(x - x_1)\]So you would plug in the values. The point you have is (4,-4). Which is x1 and which is y1?

OpenStudy (anonymous):

@Jazy I keep getting 0

OpenStudy (anonymous):

Oh and "m" is the slope.

OpenStudy (anonymous):

You shouldn't be getting a number as your answer. Are you plugging in the values?

OpenStudy (anonymous):

x1 = 4 y1 = -4 m = 4 y - (-4)= (4)(x-4) Does that make sense?

OpenStudy (anonymous):

@jazy that makes sense. so idk why i keep ending up with zero